# Multi-Shot Character Consistency for Text-to-Video Generation

### v1

@@ 3.1 Notations
3.2 Training-Free Consistent Text-to-Image generation.
@@
Our work extends ConsiStory (Tewel et al., 2024) from image to video generation, addressing the interplay between motion dynamics and identity. ConsiStory operates in three steps. (1) Subject-Driven localization with extended Self-Attention (SDSA) – localizes the subject across a set of noisy generated images by aggregating cross-attention maps across layers and timesteps. To ensure subject consistency, SDSA enables each image to attend to patches of the main subject present in other image frames. This is done by extending the self-attention mechanism, allowing it to share K, V features of the subject between multiple images. Unfortunately, SDSA alone diminishes layout diversity in the generated images. Therefore, (2) Layout Diversity – reinforces diversity through two techniques: First, it incorporates Q features from a vanilla, non-consistent sampling step. Second, it applies an inference-time dropout to the shared K, V features. Finally, (3) Refinement Injection – improves consistency in finer details by injecting the O features between corresponding subject patches.
To reduce computational complexity, and enable reusable subjects, ConsiStory uses ”anchor images”: Non-anchor images observe features from anchors during generation, but not vice versa.
3.3 Flow-based feature injection
Our approach draws inspiration from TokenFlow Geyer et al. (2023), a method for text-guided video editing. TokenFlow enforces temporal consistency in the diffusion feature space by propagating features across frames based on inter-frame correspondences. For each feature at a given location in frame $f$, TokenFlow identifies similar features in two nearby keyframes and creates a new feature by blending them according to the frame’s relative position. This preserves the overall motion pattern from the input video being edited, while incorporating the text-guided modifications.
4 Method
Our goal is to generate multiple video shots that consistently portray the same character across different scenarios described by text prompts. Building on ConsiStory, it addresses the unique challenges inherent in video generation, particularly the preservation of motion dynamics alongside character consistency. Our method comprises three key components: Framewise SDSA , a novel Query injection for motion guidance , and a Deartifacted Refinement Injection . These components work together to generate multiple video shots with consistent character and ensure motion fidelity. The pipeline is illustrated in Fig. 2.
Vanilla Caching: Before initiating our main pipeline, we generate a set of video shots using a vanilla pretrained T2V model and cache the query features , $Q_{v}$, of each denoising step.
@@ 4 Method
4.1 Framewise Subject-Driven Self-Attention
Our first step builds on the Subject-Driven Self-Attention (SDSA) mechanism (Sec.  3.2) to incorporate subject features across multiple video shots by extending the self-attention mechanism. We identified two critical challenges when adapting SDSA to video generation: (1) reliably localizing the subject during video denoising, and (2) ensuring motion fluidity is not compromised.
We term this component Framewise-SDSA. Further technical details, including the mask estimation process and the formal definition of Framewise-SDSA, are provided in Appendix A.4. .
4.2 Two-Phase Query Injection for Motion Preservation
Prior work in ConsiStory (Sec.  3.2) demonstrated success in maintaining layout diversity for image generation through SDSA dropout and query injection. However, our experiments show that directly extending this approach to video generation produces poor results, with significant visual artifacts and compromised consistency between shots (Fig. 5) . This likely occurs because ConsiStory’s vanilla-network queries are derived from latents that are influenced by consistency-preserving mechanisms in earlier steps, rather than following an independent denoising trajectory.
Phase 2: Flow-based Consistency Integration. As denoising progresses (beyond $t_{\text{pres}}$), subject consistency becomes increasingly important. To address this, we introduce Q Flow, a technique inspired by TokenFlow Geyer et al. (2023), where flow-based query features ($Q_{f}$) are injected to incorporate subject-consistent information while preserving the original motion. Similar to Sec.  3.3, in this phase, we derive a flow map from vanilla-generated keyframes ($Q_{v}$), which provides the motion structure. We then blend subject-consistent query features ($Q_{c}$) from nearby frames, as dictated by the flow. This blending process produces $Q_{f}$, that adhere to the original motion patterns while maintaining subject consistency across frames.
By following this approach, we maintain the natural flow of motion established in Phase 1 and progressively integrate subject-consistent features without sacrificing motion quality. The formal definition of our flow-based query injection process is provided in Appendix A.5.
4.3 Refinement Feature Injection for Enhanced Consistency
4.4 Implementation details
Anchor Videos: Similar to ConsiStory, we use two anchor videos that share all features between themselves. Further implementation details are provided in Appendix A.7
5 Experiments
@@ 5.2 Qualitative Results
In the appendix, we present additional capabilities of our method. When applied to Turbo-V2 (Fig. 8), our method enables subject consistency while leveraging Turbo-V2’s enhanced motion capabilities. Fig. 9 highlights our ability to handle general subject categories , such as “woman”. Fig. 10 demonstrates the ability to render multiple subjects consistently in the same scene.
5.2.1 Ablation study
We conducted an ablation study to investigate the effects of self-attention query (Q) tokens on motion and identity, and to highlight the contributions of our method’s components.
Fig. 4 illustrates typical generations for different interventions on Q tokens when combined with the extended self-attention mechanism of Fig. 4.1. When we do not intervene in the Q tokens (Fig. 4, 4th row - “No Q Intervention”), subject identity is well-maintained across video shots, but motion quality significantly degrades. This manifests in: 1) Motion synchronization: movements become synchronized across video shots, e.g., the dog’s head turning simultaneously in all shots. 2) Reduced variability in motion style and pose: similar actions are repeated across shots, e.g., the dog’s leap, the Muppet’s centered swaying, the camera movement becomes static in the skating Muppet shot. 3) Motion artifacts: to reconcile the reduced motion variability with each scene’s text prompt, videos tend towards motion-artifacts. For example, the skating Muppet’s body appears frozen while its legs are displaced to visually accommodate the “skating” action. In contrast, combining extended attention with injected Q tokens cached from vanilla diffusion-sampled video shots (Fig. 4, 3rd row - “Full Q Preservation”) restores motion but largely loses subject identity. For instance, the Muppet’s colors revert to those of the vanilla model.
These observations reveal the dual nature of Q tokens, which is central to the tokens’ role in the generation process. Injecting vanilla Q tokens restores motion, showing their influence on movement. Simultaneously, it leads to a loss of subject identity (e.g., the Muppet’s color change), indicating that Q tokens also carry identity information.
Our approach (Fig. 4, 1st row - “Ours”) achieves a balance between both worlds. By intervening in the Q tokens throughout the generation process, we restore most of the original motion, including nuanced details like body and face orientations, postures, and natural movement of specific body parts (dog’s ears, Muppet’s hands and legs). Even the parallax style of video shots is preserved (right Muppet video shot). Our method’s effectiveness stems from our two-step process. First, Q preservation in early denoising steps establishes the motion structure before the identity is fully set. Then, the flow-based Q injection allows the Q values to evolve to better match the novel (more consistent) generation, while enforcing some structural alignment through the flow process.
Refer to caption     Refer to caption
Figure 4: Ablation Study - comparing Q token intervention strategies. (click-to-view-online) “Ours” (top row) balances character consistency and natural motion. VideoCrafter2 (second row) shows diverse motion but inconsistent characters. “Full Q Preservation” (third row) directly injects Q tokens from the vanilla model without flow-based processing, preserving original motion but losing character consistency. “No Q Intervention” (bottom row) maintains strong character consistency but suffers from motion degradation and synchronization across shots.
@@ 5.3 Quantitative evaluation
Benchmark Dataset: We constructed a benchmark dataset of 30 video sets, each containing 5 video-shots with shared subjects but varying prompts. See further details in Appendix A.6.
@@
Evaluation Metrics: Our evaluation approach builds on previous work in image consistency and personalization Tewel et al. (2024); Gal et al. (2022); Ruiz et al. (2022), focusing on multi-shot set-consistency and motion dynamics. For set-consistency, we measure average pairwise DINO feature similarity Caron et al. (2021); Huang et al. (2024) across all frames in a set, excluding pairs within the same video shot. We isolate the subject by masking out the background Fu et al. (2023) before extracting each frame’s features, using ClipSEG Lüddecke & Ecker (2021) with a dynamic threshold determined by “Otsu’s method” Otsu (1979). For motion dynamics, we evaluate all $150$ generated videos using VBench’s ”Dynamic Degree” metric Huang et al. (2024), which classifies the significance of video motion by measuring RAFT-based optical flow intensity. We focused on motion dynamics over text prompt alignment due to two challenges: actions are often visible even in videos with minimal motion, making it difficult for temporal CLIP-like models Wang et al. (2024) to distinguish between our method and baselines; also, sharing seeds across baselines lead to similar visual structures, with main differences in motion quality. We include text-similarity metrics in Table 1 (Appendix), measuring temporal CLIP similarity between each video shot and its prompt.
@@
6 Conclusion and Limitations
In this work, we introduced Video Storyboarding, a novel training-free approach for generating multi-shot video sequences with consistent characters while preserving motion quality. Overall, our method provides a significant step forward in generating coherent multi-shot video sequences, offering a practical solution to the challenge of maintaining character consistency without sacrificing motion quality.
Limitations: Our approach was developed for current open text-to-video models which only generate very short videos. It is not known how it will operate with many-second-long videos. Also, balancing identity preservation and motion quality is still not perfect. We find that Q injection may be too strong and still hurt identity. To manage this, motion preservation can be compromised by partially dropping out Q injection (See Section A.2).
Reproducibility Statement
Appendix A.7 outlines the details of our implementation, including our technical solution for fitting large batches of video shots within the available GPU memory.
Appendix A.8 provides the exact instructions given to users in the study, along with examples.
The supplemental zip files contain a ”prompts” folder, which includes two files: benchmark_prompts.yaml, featuring prompts used for experiments with automatic metrics and user studies, and qualitative_prompts.yaml, containing prompts for qualitative comparisons and the ablation study.
Appendix A.4 and A.5 describe the mathematical formulation of the framewise-SDSA and Q flow injection components.
Appendix A Appendix
A.1 Additional Results
@@
Fig. 9 demonstrates Video Storyboarding’s ability to handle general subject categories. The figure shows examples of successfully generating consistent videos for broad subject types like ”woman” and ”rabbit”, indicating the model can work effectively with superclass-level prompts rather than just specific instances. Here we kept the scene style and actions the same as in our other qualitative results, and just changed the subject to a avoid detailed description
Fig. 10 showcases Video Storyboarding’s capability to handle multiple subjects. By incorporating two subjects in the prompt of the zero-shot mask, our approach can consistently render multiple characters in the same scene, as demonstrated by examples with girl-owl and boy-teddy bear pairs.
@@
Refer to caption     Refer to caption
Figure 9: General subjects: (click-to-view-online) Video Storyboarding can successfully generate consistent subjects when given with general (superclass) subject prompts like woman, or rabbit
Refer to caption     Refer to caption
Figure 10: Multi Subject: (click-to-view-online) By prompting the zero-shot mask with two subjects, our Video Storyboarding can render two consistent subjects in a scene.
Refer to caption     Refer to caption
Refer to caption     Refer to caption Refer to caption     Refer to caption
@@
A.2 Q dropout
A.3 Self-Attention in T2V models
Let $x_{i}\in\mathbb{R}^{P\times d}$ represent the input features for frame $i$, where $P$ is the number of patches and $d$ is the feature dimension. In self-attention, each patch generates three key components: $Q_{i}$ (Query features) to search for relevant information from other patches, $K_{i}$ (Key features) to match against queries, and $V_{i}$ (Value features) containing information to aggregate. The attention map is computed as $A_{i}=\textit{softmax}(Q_{i}K_{i}^{\top}/\sqrt{d_{k}})$, and is used to combine $V_{i}$ features to produce $O_{i}$ the final Output features: $O_{i}={\bm{W}}_{O}\cdot(A_{i}\cdot V_{i})$, where ${\bm{W}}_{O}$ is a linear projection matrix. Our method intervenes in this self-attention mechanism by allowing video-frames in a generated batch of videos to attend to each other and be influenced by each other’s activations.
A.4 Framewise Subject-Driven Self-Attention - Implementation Details
@@
In video generation, subject localization becomes particularly challenging during early denoising steps, where the noise is most prominent. aggregation method proposed in ConsiStory (Sec. 3.2) proved insufficient in this context, particularly during the earliest denoising steps, leading unreliable masks both in terms of accuracy and false positive localization.
@@
A.5 Flow-based Q components injection - Formal Definition
@@
A.6 Benchmark Dataset Construction:
@@ A.7 Implementation Details
A.8 User Study Protocol

### v2

@@ B.1 Notations
B.2 ConsiStory details
@@
ConsiStory Tewel et al. (2024) operates in three steps. (1) Subject-Driven localization with extended Self-Attention (SDSA) – localizes the subject across a set of noisy generated images by aggregating cross-attention maps across layers and timesteps. To ensure subject consistency, SDSA enables each image to attend to patches of the main subject present in other image frames. This is done by extending the self-attention mechanism, allowing it to share K, V features of the subject between multiple images. Unfortunately, SDSA alone diminishes layout diversity in the generated images. Therefore, (2) Layout Diversity – reinforces diversity through two techniques: First, it incorporates Q features from a vanilla, non-consistent sampling step. Second, it applies an inference-time dropout to the shared K, V features. Finally, (3) Refinement Injection – improves consistency in finer details by injecting the O features between corresponding subject patches.
Figure 8: Video Storyboarding Architecture: Our consistent denoising process has two phases: Q Preservation and Q Flow. We first generate and cache video shots using “vanilla” VideoCrafter2. In Q Preservation ($T\rightarrow t_{pres}$), we use Vanilla Q Injection to maintain motion structure by replacing our Q values with vanilla ones. In Q Flow ($t_{pres}\rightarrow t_{0}$), we use a flow map from vanilla key frames to guide Q feature injection. This phase maintains character identity by allowing the use of Q features from our consistent denoising process, while the flow map ensures that these identity-preserving features are applied in a way that’s consistent with the original motion. Throughout, we employ two complementary techniques: framewise subject-driven self-attention for visual coherence, and refinement feature injection (Section B.4) to reinforce character consistency across diverse prompts.
The pipeline is illustrated in Fig. 8.
B.3 Framewise Subject-Driven Self-Attention
Our first step builds on the Subject-Driven Self-Attention (SDSA) mechanism Tewel et al. (2024) to incorporate subject features across multiple video shots by extending the self-attention mechanism. We identified two critical challenges when adapting SDSA to video generation: (1) reliably localizing the subject during video denoising, and (2) ensuring motion fluidity is not compromised.
We term this component Framewise-SDSA. Further technical details, including the mask estimation process and the formal definition of Framewise-SDSA, are provided in Appendix B.11..
Prior work in ConsiStory (Sec.  B.2) demonstrated success in maintaining layout diversity for image generation through SDSA dropout and query injection. However, our experiments show that directly extending this approach to video generation produces poor results, with significant visual artifacts and compromised consistency between shots (Fig. 10). This likely occurs because (1) Consistory’s query injection is applied for shorter periods compared to the amount required in video models, and (2) since ConsiStory’s vanilla-network queries are derived from latents that are influenced by consistency-preserving mechanisms in earlier steps, rather than following an independent denoising trajectory.
Phase 2: Flow-based Consistency Integration. As denoising progresses (beyond $t_{\text{pres}}$), subject consistency becomes increasingly important. To address this, we introduce Q Flow, a technique inspired by TokenFlow Geyer et al. (2023), where flow-based query features ($Q_{f}$) are injected to incorporate subject-consistent information while preserving the original motion. Similar to Geyer et al. (2023), in this phase, we derive a flow map from vanilla-generated keyframes ($Q_{v}$), which provides the motion structure. We then blend subject-consistent query features ($Q_{c}$) from nearby frames, as dictated by the flow. This blending process produces $Q_{f}$, that adhere to the original motion patterns while maintaining subject consistency across frames.
By following this approach, we maintain the natural flow of motion established in Phase 1 and progressively integrate subject-consistent features without sacrificing motion quality. The formal definition of our flow-based query injection process is provided in Appendix B.12.
B.4 Refinement Feature Injection for Enhanced Consistency
B.5 Implementation details
Anchor Videos: Similar to ConsiStory, we use two anchor videos that share all features between themselves. Further implementation details are provided in Appendix B.15
B.6 Consistent Video Generation - Comparisons to Baselines
@@ B.8 Qualitative Results
When applied to Turbo-V2 (Fig. 13), our method enables subject consistency while leveraging Turbo-V2’s enhanced motion capabilities.
@@ B.9 Quantitative evaluation
Benchmark Dataset: We constructed a benchmark dataset of 30 video sets, each containing 5 video-shots with shared subjects but varying prompts. See further details in Appendix B.13.
@@
Evaluation Metrics: Our evaluation approach builds on previous work in image consistency and personalization Tewel et al. (2024); Gal et al. (2022); Ruiz et al. (2022), focusing on multi-shot set-consistency and motion dynamics. For set-consistency, we measure average pairwise DINO feature similarity Caron et al. (2021); Huang et al. (2024) across all frames in a set, excluding pairs within the same video shot. We isolate the subject by masking out the background Fu et al. (2023) before extracting each frame’s features, using ClipSEG Lüddecke & Ecker (2021) with a dynamic threshold determined by “Otsu’s method” Otsu (1979). For motion dynamics, we evaluate all $150$ generated videos using VBench’s ”Dynamic Degree” metric Huang et al. (2024), which classifies the significance of video motion by measuring RAFT-based optical flow intensity. We focused on motion dynamics over text prompt alignment due to two challenges: actions are often visible even in videos with minimal motion, making it difficult for temporal CLIP-like models Wang et al. (2024b) to distinguish between our method and baselines; also, sharing seeds across baselines lead to similar visual structures, with main differences in motion quality. We include text-similarity metrics in Table 3 (Appendix), measuring temporal CLIP similarity between each video shot and its prompt.
@@
B.10 Additional Results
@@
Refer to caption     Refer to caption Refer to caption     Refer to caption Refer to caption     Refer to caption
@@
B.11 Framewise Subject-Driven Self-Attention - Implementation Details
@@
In video generation, subject localization becomes particularly challenging during early denoising steps, where the noise is most prominent. aggregation method proposed in ConsiStory (Sec. B.2) proved insufficient in this context, particularly during the earliest denoising steps, leading to unreliable masks both in terms of accuracy and false positive localization.
@@
B.12 Flow-based Q components injection - Formal Definition
@@
B.13 Benchmark Dataset Construction:
@@ B.15 Implementation Details
Denoising Process: Videos were sampled the default VideoCrafter2 configuration, using 50 DDPM steps with a guidance scale of 12.
@@
B.16 User Study Protocol